4. The marketing manager of the Summers Company must formulate a recommendation concerning the price to be charged for a new product. According to the best available estimates, the marginal cost of the new product will be Rs. 18, and the price elasticity of demand for this product will be 3.0. (a) What recommendation should she make, if Summers wants to maximize profit? (b) If her recommendation is accepted, what will be the new product's marginal revenue?


1
Expert's answer
2022-02-24T08:16:00-0500

(a) What recommendation should she make, if Summers wants to maximize profit?

Given that the new product's price elasticity of demand is 3.0, it's clear that demand for it is very elastic. To put it another way, the quantity demanded will fall more than the amount of price increase, and vice versa.

Marginal revenue must equal marginal cost in order to maximize profit.

To get the price he will fix to get maximum profit, we use the following formula:


P = MC("\frac{n}{n+1}")

where:

P is the price of the new product

MC is the marginal cost of the new product.

n is the price elasticity of demand for the new product


P = 18("\frac{3}{3+1}") = 13.5

Therefore, Price = Rs. 13.5

Thus, to get maximum profit, he needs to put Rs. 13.5 as the price of the new product.


(b) If her recommendation is accepted, what will be the new product's marginal revenue?

Marginal revenue must equal marginal cost in order to maximize profit. Therefore, the new product's marginal revenue is Rs. 18